Suppose van de graaff generator builds a negative static charge, and a grounded conductor is placed near enough to it so that a 7.0x10^-6C of a negative charge arcs to the conductor. Estimate the number of electrons involved.
Answer #1
charge of electron e = 1.6*10-19 C
                           1C = 6.25*1018 electrons.
number of electrons = (7.0x10-6)(6.25*1018 )
                               = 4.37*1013 electrons